### v2

@@ Abstract
In recent years, the rapid evolution of computer vision has seen the emergence of various foundation models, each tailored to specific data types and tasks. In this study, we explore the adaptation of these models for few-shot semantic segmentation. Specifically, we conduct a comprehensive comparative analysis of four prominent foundation models: DINO V2, Segment Anything, CLIP, Masked AutoEncoders, and of a straightforward ResNet50 pre-trained on the COCO dataset. We also include 5 adaptation methods, ranging from linear probing to fine tuning. Our findings show that DINO V2 outperforms other models by a large margin, across various datasets and adaptation methods. On the other hand, adaptation methods provide little discrepancy in the obtained results, suggesting that a simple linear probing can compete with advanced, more computationally intensive, alternatives  [^1] Code to be released at: https://github.com/RedaBensaidDS/Foundation_FewShot
Keywords:
Deep Learning Foundation models Semantic Segmentation Few-Shot Learning
@@ 1 Introduction
Semantic segmentation, the task of pixel-level object classification within images, plays a pivotal role in computer vision applications [22]. In recent years, few-shot semantic segmentation has emerged as a challenging yet essential frontier [31]. This paradigm seeks to adapt models to novel object classes with minimal labeled data. A standard pipeline for few-shot semantic segmentation typically comprises two key components: a pretrained feature extractor, often derived from a vision foundation model, and an adapter method [2]. This setup leverages the knowledge embedded within the feature extractor and adapts it to the limited labeled data without the risk of overfitting. While extensive discussions regarding the selection of feature extractors for few-shot classification have been well-documented [21], the question of the optimal feature extractor for few-shot semantic segmentation remains largely unaddressed in the literature.
CityscapesCOCOPPD0020202020404040406060606080808080100100100100mIoUDINO V2SAMResNetCLIPMAE
@@
The primary objective of our research is to investigate which vision foundation model yields the most effective few-shot semantic segmentation pipeline. We are the first, to our knowledge, to systematically test the adaptation capabilities of renowned foundation models to this specific challenge. To this end, we introduce a novel benchmark, constructed using three widely-recognized semantic segmentation datasets, and systematically evaluate multiple adaptation methods. The four foundation models under consideration are DINO V2 [27], Segment Anything (SAM) [15], CLIP [28], and a Masked AutoEncoder [12] pretrained on ImageNet-1K. Our experimental results consistently demonstrate the superior performance of DINO V2 across various datasets.
In summary, from our study we claim the following key contributions:
  • We introduce a standardized benchmark for the evaluation of few-shot semantic segmentation in the context of vision foundation models.
  • Based on this benchmark, we conduct a comprehensive comparison of five feature extractors (four foundation models) and five adaptation methods.
  • We conduct extensive experiments to better understand the elementary contribution of foundation models design (e.g. type of architecture, size, training dataset) onto the obtained performance.
@@
Our findings indicate that DINO V2 consistently outperforms all other models across multiple datasets and adaptation methods, with SAM yielding competitive results, particularly on some datasets (Figure 1). Additionally, we find that simple and efficient adaptation methods produce results comparable to more complex counterparts, showing yet again the importance of feature extraction on top of that of downstream segmentation heads. We anticipate that our research will assist in selecting appropriate solutions for few-shot semantic segmentation tasks and contribute to the rigorous benchmarking of this evolving field.
@@ 2.1 Task Formulation
In accordance with the established terminology of few-shot classification [32], our benchmark revolves around the concept of a support set $\mathcal{S}$ and a query set $\mathcal{Q}$. The support set $\mathcal{S}$ comprises a limited number of images, each accompanied by their corresponding ground truth segmentation. These support images are employed for the training or calibration of a generic semantic segmentation model. In contrast, the query set $\mathcal{Q}$ serves as the evaluation dataset on which we compute the mean Intersection-over-Union (mIoU). In practical applications, $\mathcal{Q}$ is typically not available and is utilized exclusively for benchmarking purposes.
@@ 2.2 Benchmark Generation
    Following an arbitrary order of classes, we randomly select $k$ images from $\mathcal{D}_{i}$ without replacement to be incorporated into the support set $\mathcal{S}$. We ensure that images are only selected once.
@@
    In the event that there are insufficient remaining images for a particular class $i$, we restart from Step 1. It is important to note that this reset condition did not occur in the course of our experiments.
@@ 3 Datasets
Cityscapes [3] is a large scale dataset on semantic understanding of urban scene streets. It contains 2,975 images for training, 500 for validation and 1,525 for testing. It is annotated using 19 classes (such as “car”, “road”, “building”) and has input resolution 1024x2048.
COCO (Microsoft Common Objects in Context) [19] is a large scale dataset used for multiple tasks including semantic segmentation. The 2017 release contains 118k images for training, 5k for validation and 41k for testing. It is annotated using 80 classes (such as “person”, “bicycle”, “elephant”). Images in COCO often exhibit varying input resolutions, commonly falling between 400 and 640 pixels for both width and height.
PPD (PLANT PHENOTYPING DATASETS) [23] is a small dataset used for multiple tasks including plant segmentation (foreground to background) and leaf segmentation (multi-instance segmentation). We focus in our work on the plant segmentation task, where the dataset contains 783 images and 2 classes (foreground and background). Typical input resolutions for PPD images hover around 500 pixels for both width and height. Contrary to the two other datasets, we considered the background to be a class of its own for benchmark generation, meaning that a 1-shot problem would contain 2 support images.
@@ 4 Backbones and Adaptation Methods
In this section, we introduce the methodologies that form the basis of our comparative study. Drawing from existing literature on both few-shot classification [21] and few-shot semantic segmentation [2], we have identified that the majority of methods rely on a combination of two fundamental components.
The first essential component is a pretrained model, typically a transformer or a convnet. Notably, these pretrained models are not necessarily designed specifically for semantic segmentation but offer valuable preexisting knowledge.
@@ 4.1 Pretrained Models
In our study, we investigate five pretrained models, each offering unique characteristics and advantages (see also summary in Table 13):
@@
MAE (Masked AutoEncoder) [12] is a model pretrained on ImageNet-1K [6] using a self-supervised technique involving masking portions of input images and reconstructing them using partial tokens from the unmasked segments. It comprises a Vision Transformer (ViT) [7] component that maps observed image regions into latent representations and a decoder for image reconstruction. Previous research [12] has demonstrated MAE’s effectiveness in transfer learning for segmentation tasks, particularly in semantic segmentation.
@@
SAM (Segment Anything Model) [15] is a recent foundation model for segmentation, pretrained on the SA-1B dataset, which includes 11 million images and 1 billion segmentation masks. SAM exhibits robust generalization capabilities across a broad range of segmentation tasks. It utilizes a Vision Transformer (ViT) [7] initialized with a MAE [12], a prompt encoder, and a mask decoder. SAM’s image encoder generates embeddings from input images, which are subsequently employed by the mask decoder in conjunction with the prompt encoder. While SAM’s mask decoder outputs class-agnostic masks intended for segmenting instances or objects at various granularities, our specific task focuses exclusively on the encoder’s role as a feature extractor and does not require prompts.
@@
DINO v2 [27] is a foundational vision model trained using the self-supervised technique [1] called DINO. This methods consists in having two networks with the same architecture, namely the teacher and the student. The student parameters are learned by minimizing a loss measuring the discrepancy between teacher and student outputs learning a “local to global” correspondence where the student processes global and local views of the image while the teacher only processes the global views. The teacher is built with an exponential moving average of past iteration of the student. The model comprises a ViT trained on a diverse range of datasets, including notable semantic segmentation datasets like Cityscapes [3], ADE20k [42, 43], and Pascal VOC 2012 [8]. DINO v2 has previously been evaluated on (not few-shot) semantic segmentation datasets in [27], utilizing the frozen backbone and a simple segmentation decoder.
@@
CLIP (Contrastive Language-Image Pre-training) [28] is a model that consists of two encoders: a vision encoder and a text encoder. These encoders are jointly trained using a contrastive loss to align embeddings of images with their corresponding captions. The model is pretrained on a private dataset of 400 million (image, text) pairs and has gained significant attention for its utility in zero-shot classification on downstream tasks. In our study, we exclusively leverage the ViT version of the visual encoder component of CLIP.
@@
FCN (Fully Convolutional Network) [20] is a conventional convolutional architecture extensively used for semantic segmentation tasks. It comprises an encoder and a decoder, with the former composed of convolutional layers and the latter incorporating deconvolutional layers. In our configuration, the encoder is based on ResNet50 [13] and is pretrained on a subset of COCO, specifically using the 20 classes shared with the Pascal VOC dataset. We incorporate FCN as a convolution-only benchmark for comparison.

### v3

@@ Abstract
Few-shot semantic segmentation (FSS) is a crucial challenge in computer vision, driving extensive research into a diverse range of methods, from advanced meta-learning techniques to simple transfer learning baselines. With the emergence of vision foundation models (VFM) serving as generalist feature extractors, we seek to explore the adaptation of these models for FSS. While current FSS benchmarks focus on adapting pre-trained models to new tasks with few images, they emphasize in-domain generalization, making them less suitable for VFM trained on large-scale web datasets. To address this, we propose a novel realistic benchmark with a simple and straightforward adaptation process tailored for this task. Using this benchmark, we conduct a comprehensive comparative analysis of prominent VFM and semantic segmentation models. To evaluate their effectiveness, we leverage various adaption methods, ranging from linear probing to parameter efficient fine-tuning (PEFT) and full fine-tuning. Our findings show that models designed for segmentation can be outperformed by self-supervised (SSL) models. On the other hand, while PEFT methods yields competitive performance, they provide little discrepancy in the obtained results compared to other methods, highlighting the critical role of the feature extractor in determining results. To our knowledge, this is the first study on the adaptation of VFM for FSS.  [^1] The code is available at: https://github.com/RedaBensaidDS/Foundation_FewShot.
@@ 1 Introduction
Semantic segmentation, the task of pixel-level object classification within images, plays a pivotal role in computer vision applications (Minaee et al., 2022). This precise level of image understanding facilitates numerous applications across various domains, such as autonomous driving (Ha et al., 2017), seismic imagery analysis (Civitarese et al., 2019), aerial imagery (Tang et al., 2023) and medical imaging (Petitjean et al., 2015; Wang et al., 2019) where precise image analysis is crucial.
In recent years, few-shot semantic segmentation (FSS) (Catalano & Matteucci, 2023) has emerged as a challenging yet crucial area of research (Shaban et al., 2017). This paradigm seeks to train or adapt models to previously unseen object classes with minimal labeled data. The typical benchmarks for FSS involve datasets split into four folds with different classes: three are used for training or creating training episodes, while the fourth fold is reserved for testing or generating few-shot tasks, all containing an equal number of classes (Shaban et al., 2017; Nguyen & Todorovic, 2019).
Recently, Vision Foundation Models (VFM) have emerged as a pivotal advancement in computer vision. These models are extensively pre-trained on vast amounts of data and then fine-tuned for specific downstream tasks. Unlike traditional models trained from scratch on task-specific data, VFM learn rich, general-purpose representations, which can be transferred to a wide range of tasks and domains. The key advantage of VFM lies in their ability to capture complex patterns and semantic relationships present in diverse datasets, making them highly effective for tasks requiring cross-domain generalization. This has led us to explore the potential of adapting VFMs for FSS.
Although PASCAL-$5^{i}$ and COCO-$20^{i}$ serve as standard benchmarks for FSS (Catalano & Matteucci, 2023), they exhibit several limitations and do not align well with the capabilities of VFMs, which, we argue, allow for more general, efficient, and realistic use cases. More specifically, current benchmarks have the following limitations:
  • They focus mainly on in-domain generalization, where models are trained on specific classes and then evaluated on few-shot tasks from the remaining classes within the same dataset. VFMs, by contrast, can generalize to new data without the need to observe distinct yet similar distributions from the same dataset beforehand. As such, VFMs can be deployed in more diverse conditions.
  • They assume that the classes of interest were not seen during pretraining of the VFM, which is unverifiable when it comes to foundation models, and also a restrictive setting when it comes to actual applications.
  • They only employ binary masks during training phases, where a single class is annotated as foreground and all other pixels are treated as background in a support image [^2] While existing FSS benchmarks do support multi-class segmentation in the final query images, they often rely on binary masks during the training phase for each class in the N-way K-shot setting.. This approach, adapted from few-shot classification, enables controlled construction of few-shot episodes by ensuring a fixed number of samples per class. However, it introduces inconsistencies, such as the same class being labeled as foreground in one episode and background in another, which can hinder learning (Yang et al., 2021). Additionally, it limits the use of multi-class context during training, which is often present in realistic scenes. [^3] The recent Generalized Few-shot Semantic Segmentation (GFSS) benchmark (Tian et al., 2022) supports multi-class segmentation. However, this benchmark relies on the existence of base classes, which is incompatible with our benchmark as stated above.
  • They exhibit balanced class distributions when constructing few-shot tasks. This design simplifies performance evaluation, but does not reflect the natural statistics of real-world semantic segmentation datasets, where certain classes occur far more frequently than others. As a result, benchmarks with artificially balanced class sampling may misrepresent the challenges models face in realistic deployment scenarios.
In this paper, we introduce a new realistic benchmark meant to improve the following aspects:
  • Our benchmark does not rely on training or fine-tuning on a subset of classes before evaluating on the rest. Instead, we directly perform few-shot adaptation on the target task, without prior exposure to the dataset. This design better reflects real-world deployment scenarios of VFMs, which are expected to generalize to novel tasks with minimal supervision and without task-specific pretraining.
  • A stronger emphasis on adaptation rather than requiring the model to generalize to unseen classes in few-shot tasks  [^4] Recent benchmarks have started to address cross-domain few-shot segmentation. However, these approaches often involve training or meta-training on one dataset and then testing on another, which still differs from our objective., which aligns better with the goal of VFMs (Bommasani et al., 2021, Section 4.3).
  • Our approach uses fully labeled few-shot examples, allowing for multi-class supervision. This setup enables richer contextual understanding, avoids semantic inconsistencies during training, and supports more complex few-shot scenarios. It is also compatible with standard semantic segmentation datasets and aligns with application domains, such as autonomous driving and remote sensing—where full-scene understanding is typically required (Çağrı Kaymak & Uçar, 2018; Peng et al., 2022; Holder & Shafique, 2022).
  • Class balance is not strictly enforced, instead, our benchmark reflects the natural class distribution present in the selected datasets. For instance, in datasets like Cityscapes, frequently occurring classes such as "car" or "road" appear in most images, while other classes are comparatively rare. Preserving this imbalance provides a more realistic evaluation of model performance, particularly in terms of handling dominant and rare classes under practical conditions.
CityscapesCOCOPPD0020202020404040406060606080808080100100100100mIoUDINOv2SegFormeriBOTSAMResNetCLIPMAE
@@
The primary objective of our research is to investigate which vision foundation model and adaptation method yields the most effective few-shot semantic segmentation pipeline. We are the first, to our knowledge, to systematically test the adaptation capabilities of renowned foundation models to this specific challenge. To this end, we have introduced a novel benchmark, constructed using three widely-recognized semantic segmentation datasets, and systematically evaluated multiple adaptation methods. The five foundation models under consideration are DINOv2 (Oquab et al., 2023), Segment Anything (SAM) (Kirillov et al., 2023), CLIP (Radford et al., 2021), Masked AutoEncoder (MAE) (He et al., 2021) and iBOT (Zhou et al., 2022a). Additionally, we have carried out extensive experiments to better understand the elementary contribution of foundation models design (e.g. type of architecture, size, training dataset) and the effect of the adaptation methods onto the obtained performance.
@@
Our findings indicate that, surprisingly, models designed for segmentation are consistently outperformed by a self-supervised learning (SSL) model, DINOv2, across various datasets and adaptation methods. While Segment Anything (SAM) generally provides competitive results, it underperforms on some datasets (Figure  1). Additionally, our study indicates that parameter-efficient fine-tuning (PEFT) methods yield highly competitive performance, although the performance differential with other methods is not substantial. This underscores the critical role of feature extractors in model performance. We anticipate that our research will assist in selecting appropriate solutions for few-shot semantic segmentation tasks and contribute to the rigorous benchmarking of this evolving field.
2 Related Work
In this section, we discuss related work in the literature in few-shot classification and FSS.
2.1 Few-Shot Classification
Few-shot classification initially focused on in-domain tasks, with benchmarks involving dataset splitting into training, validation, and fake few-shot task generation for evaluation (Ravi & Larochelle, 2017; Snell et al., 2017; Vinyals et al., 2016; Finn et al., 2017). The introduction of Meta-dataset (Triantafillou et al., 2020) marked a shift towards the cross-domain setting, where the dataset used for model training differs from the tasks evaluated later, aligning with more realistic scenarios. More recently, the field has embraced the use of foundation models, adapting them to few-shot tasks. Notably, benchmarks such as those in (Zhou et al., 2022b) explore the application of CLIP (Radford et al., 2021) to 11 downstream tasks, and DINO(Caron et al., 2021) to few-shot downstream tasks(Luo et al., 2023).
Methodologically, diverse approaches have emerged, ranging from hallucination-based (Li et al., 2020; Hariharan & Girshick, 2017) to meta-learning (Finn et al., 2017; Munkhdalai et al., 2018; Zhang et al., 2021; Munkhdalai & Yu, 2017). Straightforward methods have also demonstrated effectiveness (Li et al., 2021; Tian et al., 2020; Li et al., 2022). Recent surveys, such as (Luo et al., 2023), suggest that optimal approaches often involve simple finetuning atop competitive pretrained models, augmented with an additional linear layer. This observation motivates the methods introduced in our paper.
2.2 Semantic Segmentation and Few-Shot Semantic Segmentation
Few-shot semantic segmentation has been an active area of research for several years. Early works, such as (Shaban et al., 2017) and (Nguyen & Todorovic, 2019), introduced benchmarks like PASCAL-$5^{i}$ and COCO-$20^{i}$, which share similarities with the in-domain benchmarks commonly used in few-shot classification.
Subsequent studies leveraged pretrained models on ImageNet (Yang et al., 2021; Hong et al., 2022; Wu et al., 2021; Okazawa, 2022), drawing parallels to the Meta-dataset approach in few-shot classification. Despite this progression, these methods continued to rely on the earlier benchmarks, limiting their applicability to broader scenarios.
More recently, efforts have shifted toward cross-domain few-shot semantic segmentation (Wang et al., 2022; Nie et al., 2024; Lei et al., 2022). These approaches often depend on meta-training and assume that images from the target domain originate from a distribution distinct from the training dataset. However, this assumption may not hold in scenarios involving foundation models, where domain boundaries are less clearly defined.
Recent works focused on adapting vision foundation models to FSS (Shuai et al., 2023; Liu et al., 2024; Sun et al., 2024; Meng et al., 2024; Zhu et al., 2024; Zhang et al., 2024), However, these studies mainly explore specific model adaptations, without offering a broad comparative analysis of their effectiveness as feature extractors. Our work provides a benchmark and a comprehensive comparative analysis of these models alongside various adaptation methods, making our contribution orthogonal to existing research.
Additionally, many existing methods focus on binary labeling for the few-shot setting, where only a single object of interest is annotated in the support images, and all other objects are treated as background. This simplification can hinder effective training, particularly in tasks requiring more nuanced multi-class segmentation.
@@ 3.1 Task Formulation
In accordance with the established terminology of few-shot classification (Snell et al., 2017), our benchmark revolves around the concept of a support set $\mathcal{S}$ and a query set $\mathcal{Q}$. The support set $\mathcal{S}$ comprises a limited number of images, each accompanied by their corresponding, fully labeled, ground truth segmentation. These support images are employed for the training or calibration of a generic semantic segmentation model. In contrast, the query set $\mathcal{Q}$ serves as the evaluation dataset on which we compute the mean Intersection-over-Union (mIoU). In practical applications, $\mathcal{Q}$ is typically not available and is utilized exclusively for benchmarking purposes.
@@ 3.2 Benchmark Generation
    Following an arbitrary order of classes, we randomly select $k$ images from $\mathcal{D}_{i}$ without replacement to include in the support set $\mathcal{S}$. We ensure that images are only selected once.
@@
    If there are not enough remaining images for a specific class $i$, we return to Step 1. It is important to note that this reset condition did not arise during our experiments.
@@ 4 Datasets
Cityscapes (Cordts et al., 2016; 2015) is a large scale dataset on semantic understanding of urban scene streets. It contains 2,975 images for training, 500 for validation and 1,525 for testing. It is annotated using 19 classes (such as “car”, “road”, “building”) and has input resolution 1024x2048.
COCO (Microsoft Common Objects in Context) (Lin et al., 2014) is a large scale dataset used for multiple tasks including semantic segmentation. The 2017 release contains 118k images for training, 5k for validation and 41k for testing. It is annotated using 80 classes (such as “person”, “bicycle”, “elephant”). Images in COCO often exhibit varying input resolutions, commonly falling between 400 and 640 pixels for both width and height.
PPD (PLANT PHENOTYPING DATASETS) (Minervini et al., 2016; Hanno Scharr; Sotirios A. Tsaftaris, ) is a small dataset used for multiple tasks including plant segmentation (foreground to background) and leaf segmentation (multi-instance segmentation). We focus in our work on the plant segmentation task, where the dataset contains 783 images and 2 classes (foreground-background segmentation). Typical input resolutions for PPD images hover around 500x500. Contrary to the two other datasets, we considered the background to be a class of its own for benchmark generation, meaning that a 1-shot problem would contain 2 support images.
The inclusion of these three datasets ensures a diverse evaluation framework, spanning different levels of difficulty and data distributions. From simpler binary segmentation tasks, such as foreground-background segmentation in PPD, to complex multi-class problems like COCO with 80 classes, this benchmark enables a comprehensive assessment of model performance across varied scenarios. Foreground-background segmentation, exemplified by PPD, holds significant relevance in domains like medical imaging (Dumitru & Peteleaza, 2023) and road extraction (Aich et al., 2018). Despite its smaller scale, PPD demonstrates the importance of addressing practical, real-world segmentation challenges that demand robust and efficient solutions. This approach aligns with practices in few-shot classification, such as the Coop benchmark, which utilized 11 publicly available datasets (Zhou et al., 2022b).
@@ 5 Backbones and Adaptation Methods
In this section, we introduce the methodologies that form the basis of our comparative study. Drawing from existing literature on both few-shot classification (Luo et al., 2023) and few-shot semantic segmentation (Catalano & Matteucci, 2023), we have identified that the majority of methods rely on a combination of two fundamental components.
The first essential component is a pretrained model, typically a transformer (Vaswani et al., 2017) or a convnet. It’s important to note that while these models are not specifically designed for semantic segmentation, they bring a substantial amount of valuable knowledge that can be leveraged in this context.
@@ 5.1 Pretrained Models
In our study, we investigate five pretrained models, each offering unique characteristics and advantages (see also summary of the models and their main components in Table 10):
@@
MAE (Masked AutoEncoder) (He et al., 2021) is a model pretrained on ImageNet-1K (Deng et al., 2009) using a self-supervised technique involving masking portions of input images and reconstructing them using partial tokens from the unmasked segments. It comprises a Vision Transformer (ViT) (Dosovitskiy et al., 2021) component that maps observed image regions into latent representations and a decoder for image reconstruction. Previous research (He et al., 2021) has demonstrated MAE’s effectiveness in transfer learning for segmentation tasks, particularly in semantic segmentation.
@@
SAM (Segment Anything Model) (Kirillov et al., 2023) is a recent foundation model for segmentation, pretrained on the SA-1B dataset, which includes 11 million images and 1 billion segmentation masks. SAM exhibits robust generalization capabilities across a broad range of segmentation tasks. It utilizes a Vision Transformer (ViT) (Dosovitskiy et al., 2021) initialized with a MAE (He et al., 2021), a prompt encoder, and a mask decoder. SAM’s image encoder generates embeddings from input images, which are subsequently employed by the mask decoder in conjunction with the prompt encoder. While SAM’s mask decoder outputs class-agnostic masks intended for segmenting instances or objects at various granularities, our specific task focuses exclusively on the encoder’s role as a feature extractor and does not require prompts.
iBOT (Zhou et al., 2022a) is a self-supervised learning model based on both masked image modeling and knowledge distillation. It comprises a ViT trained either on ImageNet-1K (Deng et al., 2009)or on ImageNet-22k and has show great generalization capabilities on multiple downstream tasks. In our study, we use the ViT trained on Imagenet-1K since it has better downstream tasks performance in (Zhou et al., 2022a).
@@
DINOv2 (Oquab et al., 2023) is a foundational vision model trained using the self-supervised technique (Caron et al., 2021) called DINO and iBOT (Zhou et al., 2022a). This methods consists in having two networks with the same architecture, namely the teacher and the student. The student parameters are learned by minimizing a loss measuring the discrepancy between teacher and student outputs learning a “local to global” correspondence where the student processes global and local views of the image while the teacher only processes the global views. The teacher is built with an exponential moving average of past iteration of the student. The model comprises a ViT trained on a diverse range of datasets, including notable semantic segmentation datasets like Cityscapes (Cordts et al., 2016), ADE20k (Zhou et al., 2017; 2019), and Pascal VOC 2012 (Everingham et al., ). DINOv2 has previously been evaluated on (not few-shot) semantic segmentation datasets in (Oquab et al., 2023), utilizing the frozen backbone and a simple segmentation decoder.
@@
CLIP (Contrastive Language-Image Pre-training) (Radford et al., 2021) is a model that consists of a vision encoder and a text encoder. These encoders are jointly trained using a contrastive loss to align embeddings of images with their corresponding captions. The model is pretrained on a private dataset of 400 million (image, text) pairs and has gained significant attention for its utility in zero-shot classification on downstream tasks. In our study, we exclusively leverage the ViT version of the visual encoder.
@@
FCN (Fully Convolutional Network) (Long et al., 2015) is a conventional convolutional architecture extensively used for semantic segmentation tasks. It consists of an encoder and a decoder, both of which are composed of convolutional layers. In our configuration, the encoder is based on ResNet50 (He et al., 2016) and is pretrained on a subset of COCO, specifically using the 20 classes shared with the Pascal VOC dataset. We incorporate FCN as a convolution-only benchmark for comparison.
SegFormer (Xie et al., 2021) is a semantic segmentation model based on a hierarchical Transformer (Vaswani et al., 2017) encoder paired with a simple, lightweight MLP decoder. In our study, we use a SegFormer trained on ADE20K (Zhou et al., 2017) dataset, specifically utilizing only the ViT encoder. This model was selected for its recent advancements in semantic segmentation, serving as a benchmark for comparison in our study.